# Can Features for Phishing URL Detection Be Trusted Across Diverse Datasets? A Case Study with Explainable AI

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2024conference: ; Khulna, Bangladesh; December 19-21, 2024
@@ 1. Introduction
Phishing attacks come in various forms, such as deceptive emails or mobile messages attached with fraudulent website URLs, all designed to trick users into revealing sensitive information or click on to malicious attachments (phish_threat_graham, ). According to recent statistics, the United States alone had a total of around $300K$ phishing victims, with financial losses exceeding \$52 million due to these attacks (phishing_stat_M.K, ). Historically, phishing website detection relied on traditional blacklisting where various publicly available blacklists like PhishTank (phishtank_portal, ) and other private blacklists are leveraged. While these black-box detection models may achieve high accuracy, they lack transparency and explainability. Due to this shortcoming, black-box models hinder trust and adoption in practice. Moreover, the dynamic nature of phishing website data involves concept drift (menon2021concept, ), which describes a situation where the relationship between the input data and the target variable varies over time in an online supervised learning environment. Although detecting the concept drift and retraining the model with newly extracted features (8455975, ) can partially resolve the issue, the overall feature importance in different deployment scenarios with different schemes of features can still be varied and not generalized.
To bridge this gap, SHAP (SHapley Additive exPlanations), a popular explainable AI (XAI) method, can be used to interpret the individual (i.e., local explanation) and overall model predictions, which can aid in the decision-making process (NIPS2017_7062_shap, ). In this paper, we propose to leverage XAI approaches to understand the generalization of phishing URL detection features across datasets. We incorporate XAI as a means to provide insight into the model’s decision-making process, shedding light on features which are more impactful in the classification of an instance as phishing versus benign. Our primary objective is to answer the following question- “Can features’ importance for phishing URL detection be trusted across diverse datasets?”. By answering the question, we want to know if certain set of features are ubiquitous for phishing URLs detection, or if the features are closely tied to a specific dataset.
@@ 2. RELATED WORKS
ML and Deep Learning (DL) based phishing detection research has seen significant progress with the advancement of the Artificial Intelligence. In literature, there are number of studies proposed to incorporate ML for phishing URL, domain, or website detection and supporting law-enforcement take-down decisions based on URL features and webpage contents (b2, ; b3, ; zamir2020phishing, ; pritom2020_covid_malwebsites, ; pritom2022_cns_website_lawenforcement, ). Sarasjati et al. (b7, ) and Preeti et al. (b8, ) presented comparison of various ML models across multiple phishing website datasets, and the usage of multiple datasets with varying class labels allows for a more robust analysis. These studies may provide insights about the best available ML models with a given set of features, but they can not provide any insights on the interpretation or explainability of the models. Next, Rugangazi et al. (b9, ) proposed an automated phishing detection strategy that picks important features using the global feature importance method to achieve high accuracy but it is limited to a single dataset and does not discuss model interpretability to assess the features generalizability for phishing detection. In another study, Ali et al. (aljofey2020_cnn_based_phishing, ) proposed to rely on URL character sequences using character-level convolutional neural network (CNN) while Tao et al. (feng2020visualizing, ) proposed character-level recurrent neural network (RNN) for phishing detection. However, these models mostly lack transparency for detection of phishing versus benign URLs, which may be targeted by adversaries by tweaking domain names to avoid detection. Moreover, Youness et al. (mourtaji2021hybrid_cnn, ) showed CNN outperforming blacklisting, lexical, content-based, and visual & behavioral similarity methods. In literature, we also find articles where natural language processing (NLP) is used with ML to effectively detect phishing attempts (sahingoz2019_ML_for_phishing, ).
@@
In phishing research, there are a number of features explored in literature, but among them lexical features are the ones those are very easily available and used mostly (verma2017s_malicious_url, ; rao2020catchphish_url, ). Other than lexical features, we observe URL statistics, HTML code, webpage javascript, webpage text, website external links, website structure, domain ranking, SSL certificates, TLD reputation, WHOIS data, DNS, and passive DNS based features (das2024modeling_hybrid_features, ; opara2024look_url_html, ; aljofey2022effective_html_url, ). We also find pre-trained transformer-based models such as BERT for feature extraction from websites (elsadig2022intelligent_bert, ), which suffers from dataset-dependency biases.
@@
Furthermore, researchers have proposed benchmark datasets for phishing detection with combination of four categories of features based on URL, website’s contents (HTML and Javascipts), external third-party (WHOIS, Google, OpenPageRank) or visual similarity using common sources (e.g., PhishTank, OpenPhish, Alexa, PhishMonger) (chiew2018building, ; hannousse2021towards, ; zeng2020diverse, ; el2020depth, ). In our case studies, we have analyzed the features that are common within these benchmark datasets. In summary, all the existing approaches either lack transparency (i.e., interoperability) or generalizability (i.e., reproducing results with another dataset context). In this paper, we showcase how XAI can bring transparency of detection models as well as how they can be leveraged to test the generalizability of models across datasets.
@@ 3.1.1. Dataset-1 ($D_{1}$)
This dataset is collected from Vrbančič et al. (b10, ) and consists of $88,647$ instances with $58,000$ being benign and $30,647$ being phishing URLs. Additionally, there are $111$ features in this dataset taken from URL attributes and web contents. The dataset annotators considered different types of feature columns based on the whole URL, domain name, URL directory, URL file name, URL parameters, resolving URL, and third-party services.
@@ 3.1.2. Dataset-2 ($D_{2}$)
This dataset is collected from Kaggle (b11, ) that consists of $19,431$ row entries with $9,716$ being benign and $9,715$ being phishing samples. There are $85$ total features in this dataset including features extracted from the URL, the HTML content, and the web domain.
@@ 3.1.4. Correcting Data Imbalance and Optimizing Data Splitting
Dataset $D_{2}$ is already balanced for both classes, while we observe imbalance in $D_{1}$, with $88,647$ benign and $30,647$ phishing instances (ratio of $2.9:1$). This imbalance can lead to biased models, as they tend to more accurate in predicting the majority class, leading to poor generalization performance when predicting minority class. To address this issue, we adopt the Synthetic Minority Over-sampling Technique (SMOTE) (b5, ), which is a popular oversampling method that generates synthetic instances of the minority class by interpolating between existing minority instances. We apply SMOTE only on the training portion of dataset $D_{1}$. This can produce more comparative model outcomes with the already balanced $D_{2}$. In this paper, we adopt a $70:30$ split ratio for the train-test data splits. After correcting class imbalance, for some experiment scenarios, we have merged the two datasets $D_{1}$ and $D_{2}$ considering the $20$ common features (defined in table 2) and created a third dataset, $D_{merge}=D_{merge(train)}+D_{merge(test)}$. To maintain the class balance and to generate fair explanation from SHAP, in $D_{merge(train)}$, we have taken 6,800 random instances for each groups (phishing and benign) in $D_{1}$’s training portion and merged them with the $D_{2}$’s training portion to get a total of $13,570$ phishing instances and $13,631$ benign instances. The test data for the merged dataset $D_{merge}$ is the direct concatenation of the test data for $D_{1}$ and $D_{2}$. Table 1 depicts the distribution of all datasets $D_{1}$, $D_{2}$, and $D_{merge}$.
@@ 3.4. Model Evaluation in Different Experiments
Common (20) Exp-5 train on $D_{1}$, test on $D_{2}$ 51.0 59.0 51.0 36.0
Exp-6 train on $D_{2}$, test on $D_{1}$ 59.0 70.0 58.0 50.0
@@ 4.1. Limitations
Even though we shed light on the phishing detection features’ deviation of effectiveness across datasets, there remain some limitations in the present study. First, the present study does not take into consideration each feature’s instance-specific contributions (local explanation) for identifying which features are deviating in the case of an individual test instance when trained on one dataset and tested on another. Second, we rely on the features’ SHAP values when interpreting the model. While SHAP gives a single measure of feature relevance, it makes the assumption that features are independent, which may not be true in all circumstances. Third, we only used two datasets in this study and maybe further datasets can be incorporated to generalize the concepts. It could be interesting if the merged dataset training can also be validated to provide reasonably good results for another third dataset containing the same common features and characteristics for testing. Fourth, We also have not considered the datasets with visual-based features as they do not often share common features. Fifth, the common phishing versus benign classification features that we have analyzed are limited to mostly lexical features as they are more commonly found across multiple datasets. Sixth, we have not considered the deep learning methods which can be scoped into future studies.
@@ 5. CONCLUSION
The usage of Explainable AI has allowed us to establish variables which are most important in detecting phishing websites, as well as how these features differ considerably between datasets. The analysis discovered critical variables that contribute significantly to the model’s predictions in several experiment scenarios. While the model performs well when trained and evaluated on the same dataset, its effectiveness is reduced when the training and test dataset are not from the same source even though they share the same set of features with similar statistics. This experimental evidences stress the need of training on different and representative datasets to improve model generalizability in practice before claiming a phishing detection model as effective with certain features. Thus, we can evidently state that phishing URL detection model that trained on singular dataset specifically using the lexical features can not be trusted across diverse datasets even if the features have very close data distributions. We recommend using multiple diverse merged dataset as a better method to use as training for the phishing detection model and always use explainable methods for AI/ML model’s verification and trustworthiness for further decision-making on the outcome. In future, more datasets can be incorporated with more diverse feature lists. The detail implementation code and dataset are shared in the following anonymous Github repository (https://anonymous.4open.science/r/Deviation-in-Feature-Contribution-7760/) for reproducing the experiments.

### v2

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: 11th International Conference on Networking, Systems, and Security; December 19–21, 2024; Khulna, Bangladeshbooktitle: 11th International Conference on Networking, Systems, and Security (NSysS ’24), December 19–21, 2024, Khulna, Bangladeshdoi: 10.1145/3704522.3704532isbn: 979-8-4007-1158-9/24/12ccs: Security and privacy Phishing Detectionccs: Computing Methodologies Artificial Intelligence
@@ 1. Introduction
Phishing attacks come in various forms, such as deceptive emails or mobile messages attached with fraudulent website URLs, all designed to trick users into revealing sensitive information or click on to malicious attachments (Thomson, 2023). Moreover, the potential abuse of generative AI and large language models may add more stress towards defenders to cope with these attacks (Roy et al., 2024; Gupta et al., 2023; Shibli et al., 2024). According to recent statistics, the United States alone had a total of around $300K$ phishing victims, with financial losses exceeding \$52 million due to these attacks (Main, 2023). Historically, phishing website detection relied on traditional blacklisting where various publicly available blacklists like PhishTank (Group, [n. d.]) and other private blacklists are leveraged. While these black-box detection models may achieve high accuracy, they lack transparency and explainability. Due to this shortcoming, black-box models hinder trust and adoption in practice. Moreover, the dynamic nature of phishing website data involves concept drift (Menon and Gressel, 2021), which describes a situation where the relationship between the input data and the target variable varies over time in an online supervised learning environment. Although detecting the concept drift and retraining the model with newly extracted features (Tan et al., 2018) can partially resolve the issue, the overall feature importance in different deployment scenarios with different schemes of features can still be varied and not generalized.
To bridge this gap, SHAP (SHapley Additive exPlanations), a popular explainable AI (XAI) method, can be used to interpret the individual (i.e., local explanation) and overall model predictions, which can aid in the decision-making process (Lundberg and Lee, 2017). In this paper, we propose to leverage XAI approaches to understand the generalization of phishing URL detection features across datasets. We incorporate XAI as a means to provide insight into the model’s decision-making process, shedding light on features which are more impactful in the classification of an instance as phishing versus benign. Our primary objective is to answer the following question- “Can features’ importance for phishing URL detection be trusted across diverse datasets?”. By answering the question, we want to know if certain set of features are ubiquitous for phishing URLs detection, or if the features are closely tied to a specific dataset.
@@ 2. RELATED WORKS
ML and Deep Learning (DL) based phishing detection research has seen significant progress with the advancement of the Artificial Intelligence. In literature, there are number of studies proposed to incorporate ML for phishing URL, malicious domain or website detection and supporting law-enforcement take-down decisions based on URL features and webpage contents (Karim et al., 2023; Chinnasamy et al., 2022; Bouijij and Berqia, 2021; Zamir et al., 2020; Pritom et al., 2020; Pritom and Xu, 2022; Abdelhamid et al., 2017; Xu et al., 2013; Jain and Gupta, 2018; Gupta et al., 2021; Maroofi et al., 2020; Yilmaz et al., 2020). However, a common limitation across these studies is the lack of interpretability of the underlying decision-making processes of the models used, often referred to as the “black box” problem. Furthermore, there is a scarcity of research that compares the performance and feature importance across diverse datasets. Sarasjati et al. (Sarasjati et al., 2022) and Preeti et al. (Preeti and Sharma, 2023) presented comparison of various ML models across multiple phishing website datasets, and the usage of multiple datasets with varying class labels allows for a more robust analysis. These studies may provide insights about the best available ML models with a given set of features, but they can not provide any insights on the interpretation or explainability of the models. Next, Rugangazi et al. (Rugangazi and Okeyo, 2023) proposed an automated phishing detection strategy that picks important features using the global feature importance method to achieve high accuracy but it is limited to a single dataset and does not discuss model interpretability to assess the features generalizability for phishing detection. In another study, Ali et al. (Aljofey et al., 2020) proposed to rely on URL character sequences using character-level convolutional neural network (CNN) while Tao et al. (Feng and Yue, 2020) proposed character-level recurrent neural network (RNN) for phishing detection. However, these models mostly lack transparency for detection of phishing versus benign URLs, which may be targeted by adversaries by tweaking domain names to avoid detection. Moreover, Youness et al. (Mourtaji et al., 2021) showed CNN outperforming blacklisting, lexical, content-based, and visual & behavioral similarity methods. In literature, we also find articles where natural language processing (NLP) is used with ML to effectively detect phishing attempts (Sahingoz et al., 2019).
@@
In phishing research, there are a number of features explored in literature, but among them lexical features are the ones those are very easily available and used mostly (Verma and Das, 2017; Rao et al., 2020). Other than lexical features, we observe URL statistics, HTML code, webpage javascript, webpage text, website external links, website structure, domain ranking, SSL certificates, TLD reputation, WHOIS data, DNS, and passive DNS based features (Das Guptta et al., 2024; Opara et al., 2024; Aljofey et al., 2022). We also find pre-trained transformer-based models such as BERT for feature extraction from websites (Elsadig et al., 2022), which suffers from dataset-dependency biases.
@@
Furthermore, researchers have proposed benchmark datasets for phishing detection with combination of four categories of features based on URL, website’s contents (HTML and Javascipts), external third-party (WHOIS, Google, OpenPageRank) or visual similarity using common sources (e.g., PhishTank, OpenPhish, Alexa, PhishMonger) (Chiew et al., 2018; Hannousse and Yahiouche, 2021; Zeng et al., 2020; El Aassal et al., 2020). In our case studies, we have analyzed the features that are common within these benchmark datasets. In summary, all the existing approaches either lack transparency (i.e., interoperability) or generalizability (i.e., reproducing results with another dataset context). In this paper, we showcase how XAI can bring transparency of detection models as well as how they can be leveraged to test the generalizability of models across datasets.
@@ 3.1.1. Dataset-1 ($D_{1}$)
This dataset is collected from Vrbančič et al. (Vrbančič et al., 2020) and consists of $88,647$ instances with $58,000$ being benign and $30,647$ being phishing URLs. Additionally, there are $111$ features in this dataset taken from URL attributes and web contents. The dataset annotators considered different types of feature columns based on the whole URL, domain name, URL directory, URL file name, URL parameters, resolving URL, and third-party services.
@@ 3.1.2. Dataset-2 ($D_{2}$)
This dataset is collected from Kaggle (Winson, 2024) that consists of $19,431$ row entries with $9,716$ being benign and $9,715$ being phishing samples. There are $85$ total features in this dataset including features extracted from the URL, the HTML content, and the web domain.
@@ 3.1.4. Correcting Data Imbalance and Optimizing Data Splitting
Dataset $D_{2}$ is already balanced for both classes, while we observe imbalance in $D_{1}$, with $88,647$ benign and $30,647$ phishing instances (ratio of $2.9:1$). This imbalance can lead to biased models, as they tend to more accurate in predicting the majority class, leading to poor generalization performance when predicting minority class. To address this issue, we adopt the Synthetic Minority Over-sampling Technique (SMOTE) (Chawla et al., 2002), which is a popular oversampling method that generates synthetic instances of the minority class by interpolating between existing minority instances. We apply SMOTE only on the training portion of dataset $D_{1}$. This can produce more comparative model outcomes with the already balanced $D_{2}$. In this paper, we adopt a $70:30$ split ratio for the train-test data splits. After correcting class imbalance, for some experiment scenarios, we have merged the two datasets $D_{1}$ and $D_{2}$ considering the $20$ common features (defined in table 2) and created a third dataset, $D_{merge}=D_{merge(train)}+D_{merge(test)}$. To maintain the class balance and to generate fair explanation from SHAP, in $D_{merge(train)}$, we have taken 6,800 random instances for each groups (phishing and benign) in $D_{1}$’s training portion and merged them with the $D_{2}$’s training portion to get a total of $13,570$ phishing instances and $13,631$ benign instances. The test data for the merged dataset $D_{merge}$ is the direct concatenation of the test data for $D_{1}$ and $D_{2}$. Table 1 depicts the distribution of all datasets $D_{1}$, $D_{2}$, and $D_{merge}$.
@@ 3.4. Model Evaluation in Different Experiments
Common Exp-5 train on $D_{1}$, test on $D_{2}$ 51.0 59.0 51.0 36.0
(20) Exp-6 train on $D_{2}$, test on $D_{1}$ 59.0 70.0 58.0 50.0
@@ 4.1. Limitations
Even though we shed light on the phishing detection features’ deviation of effectiveness across datasets, there remain some limitations in the present study. First, the study does not take into consideration each feature’s instance-specific contributions (local explanation) for identifying which features are deviating in the case of an individual test instance when trained on one dataset and tested on another. Second, we rely on the features’ SHAP values when interpreting the model. While SHAP gives a single measure of feature relevance, it makes the assumption that features are independent, which may not be true always. Third, We also have not considered any datasets with shared visual similarity features because of scarcity of such features. Fourth, the common phishing versus benign URL classification features that we have analyzed are limited to mostly lexical features as they are more commonly found across multiple datasets. Fifth, we have not considered the deep learning methods which can be considered in future studies.
@@ 5. CONCLUSION
The usage of Explainable AI has allowed us to establish variables which are most important in detecting phishing websites, as well as how these features differ between datasets. The analysis discovered critical variables that contribute significantly to the model’s predictions in several experiment scenarios. While the model performs well when trained and evaluated on the same dataset, its effectiveness is reduced when the training and test dataset are different even though they share the same features. This experimental evidences stress the need of training on different and representative datasets to improve model generalizability in practice before claiming a phishing detection model as effective with certain features. Thus, we can evidently state that phishing URL detection model that trained on singular dataset specifically using the lexical features can not be trusted across diverse datasets even if the features have very close data distributions. We recommend using multiple diverse merged dataset as a better method to use as training for the phishing detection model and always use explainable methods for AI/ML model’s verification and trustworthiness for further decision-making on the outcome. The detail implementation code and dataset are shared in the following Github repository (https://anonymous.4open.science/r/Deviation-in-Feature-Contribution-7760/) for reproducing the experiments.